Closing the Balances in The Variance Accounts at the End of the Year

Yohan Company has the following balances in its direct materials and direct labor variance accounts at year-end:

Debit Credit
Direct Materials Price Variance $13,650   
Direct Materials Usage Variance $1,270    
Direct Labor Rate Variance 890    
Direct Labor Efficiency Variance $12,640   

Unadjusted Cost of Goods Sold equals $1,510,000, unadjusted Work in Process equals $276,000, and unadjusted Finished Goods equals $240,000.

Required:

1. Assume that the ending balances in the variance accounts are immaterial and prepare the journal entries to close them to Cost of Goods Sold. Note: Close the variances with a debit balance first. If an amount box does not require an entry, leave it blank or enter "0".

Close variances with debit balance
Close variances with credit balance

What is the adjusted balance in Cost of Goods Sold after closing out the variances?

$

2. What if any ending balance in a variance account that exceeds $9,000 is considered material? (a) Close the immaterial variance accounts to Cost of Goods Sold. (b) Prorate the largest of the labor variances among Cost of Goods Sold, Work in Process, and Finished Goods on the basis of prime costs in these accounts. (c) Prorate the largest of the material variances among Cost of Goods Sold, Work in Process, and Finished Goods on the basis of prime costs in these accounts. The prime cost in Cost of Goods Sold is $1,050,000, the prime cost in Work in Process is $165,800, and the prime cost in Finished Goods is $128,000. If an amount box does not require an entry, leave it blank or enter "0".

Note: Round all interim calculations to three decimal places, and round your final answers to the nearest dollar. Adjust credit entry for rounding to ensure debits equal credits in journal entry.

(a)
(b)
(c)

What are the adjusted balances in Work in Process, Finished Goods, and Cost of Goods Sold after closing out all variances?

Adjusted balance
Work in Process $
Finished Goods $
Cost of Goods Sold $
Answer #1
1) Cost of goods sold $        26,290
Direct materials price variance $       13,650
Direct labor efficiency variance $       12,640
(To close variances with debit balance)
Direct materials usage variance $          1,270
Direct labor rate variance $              890
Cost of goods sold $         2,160
(To close variances with credit balance)
Adjusted balance in Cost of goods sold after closing out the variances = 1510000+26290-2160 = $ 15,34,130
2) a) Direct materials usage variance $          1,270
Direct labor rate variance $              890
Cost of goods sold $         2,160
(To close variances with immaterial balance)
b) Proration of direct labor efficiency variance: Prime cost Variance prorated
Cost of goods sold 1050000 9876
Work in process 165800 1560
Finished goods 128000 1204
Total 1343800 12640
Rate for proration = 12640/1343800 = 0.9406% of prime cost
JOURNAL ENTRY:
Cost of goods sold 9876
Work in process 1560
Finished goods 1204
Direct labor efficiency variance 12640
(To prorate direct labor efficiency variance)
c) Proration of direct material price variance: Prime cost Variance prorated
Cost of goods sold 1050000 10666
Work in process 165800 1684
Finished goods 128000 1300
Total 1343800 13650
Rate for proration = 13650/1343800 = 1.0158% of prime cost
JOURNAL ENTRY:
Cost of goods sold 10666
Work in process 1684
Finished goods 1300
Direct material price variance 13650
(To prorate direct material price variance)
Adjusted balance
Work in Process (276000+1560+1684) 279244
Finished Goods (240000+1204+1300) 242504
Cost of Goods Sold (1510000-2160+9876+10666) 1528382